How do you subract with fractional exponents? 2 5/2 - 2 3/2

Answer by ewatrrr(24785)


Hi
How do you subract with fractional exponents? 2 5/2 - 2 3/2
Recommend using the improper fraction format:
9/2 - 7/2 = 2/2 = 1